# MMD-Flagger: Leveraging Maximum Mean Discrepancy to Detect Hallucinations

### v1

@@ C.2 Additional results for the LFAN-HALL Dataset
In detecting the hallucination by MMD-Flagger, the MMD distance at $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ is the principal factor of forming the U-shape in the MMD distances trajectory, as shown in Figure 1; the U-shape is formed since $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ is larger than $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto})$ values. Figure 3 depicts distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ per ground-truth annotation label. These distributions represent clear differences by the annotation labels; Comparing these distributions, $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ with the hallucination label represents the high volume in the center, while the distribution with the correct label depicts the high volume at the quite low $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ value. In case of the machine translation errors, the distribution of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ looks quite different from the distribution of the correct label. There may be room for employing MMD-Flagger for detecting the machine translations; However, further studies are required.
Figure 3: Distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ by annotation label. The horizontal axes represent values of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$, and the vertical axes are percentages. Comparing two distributions “Correct” and “Hallucination”, we observe frequent larger values of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ in “Hallucination”. This tendency suggests that the U-shape of the trajectory is likely to form.
@@
Figure 4: Distributions of $\arg\min(\mathcal{T}_{\widehat{MMD}})$. The horizontal axes represent $\tau$, and the vertical axes are the percentages. Comparing two distributions “Correct” and “Hallucination”, the high frequency of being $\arg\min(\mathcal{T}_{\widehat{MMD}})$ is at $0.1$ in the “Correct” label, while $>0.1$ in the “Hallucination” label. This pattern suggests that the U-shape is likely to form in the “Hallucination” label.
Detection Performance per Label. Table 7 represents the hallucination detection performance per annotation label. The kernel settings of the MMD-Flagger are the same as described in Section 3.2.
@@
Threshold Selection. Figure 5 represents the uncertainty score distributions and Receiver operating characteristic (ROC) curves of Seq-Logprob and MC-DSim.
@@
Appendix D Datasets
The dataset description is found at Section 3.1. In this section, we describe additional information about the datasets and label statistics.
D.1 LFAN-HALL Dataset [27]
The original dataset file is provided by the authors [^8] https://github.com/deep-spin/hallucinations-in-nmt. As we discuss in Appendix A, we reviewed the annotation labels and modified it. In the following section, we show label statistics of the modified dataset.
[…]
Label Count Ratio
no-error 2,885 0.8455
hallucination 291 0.0852
mt-error 233 0.0682
hallucination,mt-error 3 0.0008
Table 9: Label distribution of the LFAN-HALL dataset. The comma “,” represents the two labels are given.
Label Count Ratio
no-error 2,885 0.8455
error-omission 193 0.0565
error-full 121 0.0354
error-strong 83 0.0243
error-repetitions, error-strong 47 0.0137
error-named-entities 37 0.0108
error-repetitions 29 0.0084
error-full, error-repetitions 11 0.0032
error-named-entities, error-omission 3 0.0008
error-omission, error-strong 1 0.0002
error-omission, error-repetitions 1 0.0002
error full, error-named-entities 1 0.0002
D.2 Halomi Dataset [11]
The label statistics of the Halomi dataset are shown in Table 10.
Table 10: Hallucination label distribution of the Halomi dataset.
Class label Count Ratio
1 No hallucination 2,131 0.74
3 Partial hallucination 291 0.10
2 Small hallucination 225 0.07
4 Full hallucination 218 0.07
Appendix E Qualitative results
The detection examples are all with the MMD-Flagger with the vector representation of avg and the kernel setting of gaussian using the $25$-th percentile value for computing the Gaussian kernel length scale, reported in Table 2.
E.1 Correct hallucination detections (True Positive)
$y_{\text{beam}}$ The room was very small and the bathroom was very small.
$y_{\text{sto}}^{\tau=0.1}$ The hotel is located in a quiet area of the city.
$y_{\text{beam}}$ Old and young people will find themselves together in the village’s respective squares for fun and fun.
$y_{\text{beam}}$ The hotel is located in the heart of the city, just a few steps away from the train station.
E.2 Correct translations (True Negative)
Reference The president concluded the debate as follows:
$y_{\text{beam}}$ At the end of the debate, the President said:
Source Politiken, auf die besonderes Augenmerk zu richten ist, und
Reference policies which should be given special attention, and
$y_{\text{beam}}$ policies to which particular attention needs to be paid; and
E.3 Detection by error (False Positive)
False Positive cases are often observed when $y_{\text{beam}}$ and $y_{\text{sto}}^{\tau=0.1}$ are semantically similar; however, dissimilarity due to different expressions or words. These false-positive errors would be due to the vector representations avg of the word embedding vectors. Due to the aggregation, the Gaussian kernel failed to compute the similarity correctly, and thus, $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ becomes a high value.
Source Ich komme zum Schluß noch darauf zu sprechen.
Reference I shall come back to this point again at the end.
$y_{\text{beam}}$ Finally, I would like to mention this.
$y_{\text{sto}}^{\tau=0.1}$ I will end by mentioning this.
Source Wollen sie eine Firma gründen oder umstrukturieren?
Reference Do you intend to market new product or service?
$y_{\text{beam}}$ Do they want to set up or restructure a company?
$y_{\text{sto}}^{\tau=0.1}$ Do you want to start a company or restructure it?
E.4 Fail to detect (False Negative)
The false-negative errors are often observed when the stochastic sampled translations $y_{\text{sto}}^{\tau=0.1}$ are really similar to the beam search translation $y_{\text{beam}}$. In this case, $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ becomes a small value and is hard to form a U-shape; therefore, MMD-Flagger failed to detect the hallucination.
Source Februar 1997 angenommenen Entschließung ABl.
Reference 61. adopted on 19 February 1997;
$y_{\text{beam}}$ Resolution OJ L 208, p.
$y_{\text{sto}}^{\tau=0.1}$ Resolution OJ L 208, p.
Source Oktober zu verlängern.
Reference police and penitentiary.
$y_{\text{beam}}$ The Commission has decided to extend the programme.
$y_{\text{beam}}$ Mr Martin BARTENSTEINBundesminister für Wirtschaft und Arbeit.
Appendix F Comparison of Kernel Length Scale Selection
The “median heuristic” [23] is a widely used method in the kernel literature for selecting the length scale of a kernel function. It computes the pairwise distance matrix (e.g., using the Euclidean distance) over a set of data points and uses the median of these distances as the length scale. While the $50$-th percentile (i.e., the median) is commonly used, other percentiles can also be considered.
@@ Appendix F Comparison of Kernel Length Scale Selection
Appendix G Hallucination Detection using Vectors from Intermediate Layers
In this section, we report MMD-Flagger detection scores using vectors at intermediate layers of a model. Previous work reports that the model’s intermediate hidden states correlate with the existence of hallucinated token sequences [11, 2, 49]. Thus, vector representations from the intermediate vector values of the outputs could also contain valuable information for the hallucination detection task.
In Section 2.2, we introduce the vector representation using the word embedding matrix $E\in\mathbb{R}^{V\times D}$. Instead of employing the matrix $E$, we access an intermediate layer of the LLM’s decoder and extract vector representations after the “Add & Norm” operation at each layer, as defined in the Transformer architecture [53], where “Add & Norm” means the residual connection followed by layer normalization. In this work, the model for the LFAN-HALL dataset has $6$ layers (the dimension size of $512$), and the model for the Halomi dataset has $12$ layers (the dimension size of $1,024$).
@@ Appendix G Hallucination Detection using Vectors from Intermediate Layers
Appendix H Detection Stability of MMD-Hallucination-Flagger
In this section, we discuss the stability of MMD-Flagger in detecting hallucinations, focusing on the impact of randomness and sample size. The detection process involves stochastic sampling, where a random selection of the next token is made based on the probability distribution generated by the softmax function, as detailed in Eq.(1). Our empirical findings suggest that MMD-Flagger is effective in identifying hallucinations with sequences comprising just $25$ tokens.
@@ Appendix H Detection Stability of MMD-Hallucination-Flagger
Figure 7 illustrates the MMD trajectory, denoted as $\mathcal{T}_{\widehat{MMD}}$, based on $10$ repetitions with varying sample sizes of $\{10,25,50,100\}$.
@@
The MMD distance $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ plays a significant role in shaping the U-curve of $\mathcal{T}_{\widehat{MMD}}$. Therefore, we analyze the distribution of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ across 294 examples.
Our assumption is that if $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ is close to the converged value, the distance value does not change significantly, regardless of the random seeds used or the sample size. As a result, the distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ are expected to exhibit similar shapes.
Figure 8 illustrates the distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ at the first and third repetitions. With a sample size of 10, the two distributions appear quite different; however, at a sample size of 100, the distribution shapes seem to converge.
(a) With $10$ samples
(b) With $25$ samples
(c) With $50$ samples
(d) With $100$ samples
Figure 7: Comparing $\mathcal{T}_{\widehat{MMD}}$ of sample sizes $\{10,25,50,100\}$. The line is the average value of $10$ repetitions, and the area plots are the variance values.
@@
Figure 8: Comparisons of distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ over $294$ examples. The left column shows the distributions at first repetition, and the right column is for the one of the third repetition.
Appendix I MMD trajectory filtering
This section introduces the hallucination detection performance after applying the trajectory filtering to $\mathcal{T}_{\widehat{MMD}}$. However, we do not observe substantial differences by applying trajectory filters.
The MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ can exhibit irregular shapes, as illustrated in Figure 9 and filtered trajectories by the average window. To address these irregularities, a smoothing method can be applied to create a more uniform trajectory. By utilizing smoothing techniques, the MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ becomes easier to manage, which may enhance the ability of MMD-Flagger to detect hallucinations more effectively.
Figure 9: An example of the MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ and filtered trajectories with the average window by $2$ and $3$ sizes. The blue solid line is the original MMD trajectory. The purple dashed line is the filtered line with the window size of $2$, and the green dashed line is the filtered line with the window size of $3$. The red circle is the minimum value of the original MMD trajectory, and black “x” marks are the minimum values on the filtered trajectory.
Table 13 compares the hallucination detection performance of the original MMD trajectory and filtered trajectories. The length scale computation at $25$-th percentile; same as Section 3.2.
Table 13: Comparing the hallucination detection of the original MMD trajectory and filtered trajectories. The score formats are “recall (precision)”.
Vector representation Kernel Filtering Window size LFAN-HALL
avg dot No Filter 0.820 (0.125)

### v3

@@ C.2 Additional results for the LFAN-HALL Dataset
In detecting the hallucination by MMD-Flagger, the MMD distance at $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ is the principal factor of forming the U-shape in the MMD distances trajectory, as shown in Figure 1; the U-shape is formed since $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ is larger than $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto})$ values. Figure 3 depicts distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ per ground-truth annotation label. These distributions represent clear differences by the annotation labels; Comparing these distributions, $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ with the hallucination label represents the high volume in the center, while the distribution with the correct label depicts the high volume at the quite low $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ value. In case of the machine translation errors, the distribution of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ looks quite different from the distribution of the correct label. There may be room for employing MMD-Flagger for detecting the machine translations; However, further studies are required.
Figure 3: Distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ by annotation label. The horizontal axes represent values of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$, and the vertical axes are percentages. Comparing two distributions “Correct” and “Hallucination”, we observe frequent larger values of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ in “Hallucination”. This tendency suggests that the U-shape of the trajectory is likely to form at the case of “Hallucination”.
@@
[…]
Detection Performance per Label.
Table 9 represents the hallucination detection performance per annotation label. The kernel settings of the MMD-Flagger are the same as described in Section 3.2.
@@
Threshold Selection
Figure 5 represents the uncertainty score distributions and Receiver operating characteristic (ROC) curves of Seq-Logprob and MC-DSim.
@@
Appendix D Evaluation using various kernel functions
D.1 Comparison of Kernel Length Scale Selection
The “median heuristic” (Garreau et al., 2017) is a widely used method in the kernel literature for selecting the length scale of a kernel function. It computes the pairwise distance matrix (e.g., using the Euclidean distance) over a set of data points and uses the median of these distances as the length scale. While the $50$-th percentile (i.e., the median) is commonly used, other percentiles can also be considered.
@@ D.1 Comparison of Kernel Length Scale Selection
D.2 Comparison of various kernel functions
We detail the five kernel functions (Schölkopf and Smola, 2002) employed for the MMD estimate, along with the specific hyperparameter ranges explored in our study. In all definitions, $a$ and $b$ denote the input vectors $\mathbb{R}^{D}$.
The Square Exponential Kernel is defined as:
[…]
where $\gamma>0$ is the inverse length-scale parameter. We tested $\gamma\in\{0.001,0.01,0.1,1.0,10\}$.
The Polynomial Kernel is defined as:
$k(a,b;d,\gamma)=\left(\gamma a^{\intercal}b+r\right)^{d},$
where $d$ is the polynomial degree. We tested $d\in\{1,2,3,4\}$ and $r$ (the intercept) $\{0,5,10\}$.
The Laplace Kernel is defined as:
[…]
where $\sigma>0$ is the length-scale parameter. We tested $\sigma\in\{0.1,0.35,0.6,...,2.0\}$.
The Sigmoid Kernel is defined as:
$k(a,b;\alpha,c)=\text{tahn}\left(\alpha a^{\intercal}b+c\right),$
where $\alpha$ is a scaling factor and $c$ is the intercept. We tested $\alpha\in\{0.01,0.1,1.0\}$ and $c\in\{-1.0,0,1.0\}$.
The Matérn kernel (Rasmussen and Williams, 2006) is defined as:
$\begin{split}k_{\nu}(a,b)&=\frac{1}{\Gamma(\nu)2^{\nu-1}}\left(\frac{\sqrt{2\nu}}{\ell}d(a,b)\right)^{\nu}\\ &K_{\nu}\left(\frac{\sqrt{2\nu}}{\ell}d(a,b)\right),\end{split}$
[…]
$k_{0.5}(r)$ $=\sigma^{2}\exp\left(-\frac{r}{\ell}\right)$
$k_{1.5}(r)$ $=\sigma^{2}\left(1+\frac{\sqrt{3}r}{\ell}\right)\exp\left(-\frac{\sqrt{3}r}{\ell}\right)$
$k_{2.5}(r)$ $=\sigma^{2}\left(1+\frac{\sqrt{5}r}{\ell}+\frac{5r^{2}}{3\ell^{2}}\right)\exp\left(-\frac{\sqrt{5}r}{\ell}\right),$
[…]
Dataset. We use a subset of the Halomi dataset (Dale et al., 2023). The subsets are two translation pairs from eng-Latn to deu-Latn ($146$ examples) and from deu-Latn to eng-Latn ($155$ examples).
The vector representation is by the word embedding layer, and vector aggregation is with the mean over vectors of $T$ tokens.
Evaluation Result. The detection result is reported at Table 11. For each kernel function, we report “recall (precision)” combination regarding the F1 score that is a harmonic mean of precision and recall. We prioritize F1 score over maximizing recall alone, as certain parameter settings achieve perfect recall (1.0) by trivially classifying all examples as hallucinations, which is not practically meaningful.
As shown in Table 11, the Polynomial kernel and the Gaussian kernel achieve the highest recall ($\mathbf{0.827}$). This indicates that a strong capability for detecting hallucinations when parameters are optimally tuned on a labeled calibration dataset. These two kernels require choosing a proper parameter on a calibration dataset having the hallucination label.
The Dot product kernel exhibits slightly inferior performance ($0.793$ recall). However, because the Dot product kernel is non-parametric and requires no calibration data, it stands as a robust and suitable baseline choice when the calibration dataset is unavailable.
Table 11: Comparison of various kernels at Appendix D.2.
Kernel Recall (Precision), parameters
Polynomial 0.827 (0.146), $d=2,r=0$
Gaussian 0.827 (0.138), $25$-th percentile
Dot 0.793 (0.133)
Square Exp. 0.517 (0.091), $\gamma=0.001$
Laplace 0.137 (0.056), $\sigma=0.1$
Matérn 0.241 (0.067), $\nu=0.5$
Sigmoid 0.275 (0.266), $\alpha=0.01,c=-1.0$
Appendix E Hallucination Detection using Vectors from Intermediate Layers
In this section, we report MMD-Flagger detection scores using vectors at intermediate layers of a model. Previous work reports that the model’s intermediate hidden states correlate with the existence of hallucinated token sequences (Dale et al., 2023; Azaria and Mitchell, 2023; Su et al., 2024). Thus, vector representations from the intermediate vector values of the outputs could also contain valuable information for the hallucination detection task.
In Section 2.2, we introduce the vector representation using the word embedding matrix $E\in\mathbb{R}^{V\times D}$. Instead of employing the matrix $E$, we access an intermediate layer of the LLM’s decoder and extract vector representations after the “Add & Norm” operation at each layer, as defined in the Transformer architecture (Vaswani et al., 2017), where “Add & Norm” means the residual connection followed by layer normalization. In this work, the model for the LFAN-HALL dataset has $6$ layers (the dimension size of $512$), and the model for the Halomi dataset has $12$ layers (the dimension size of $1,024$).
@@ Appendix E Hallucination Detection using Vectors from Intermediate Layers
Appendix F Detection Stability of MMD-Hallucination-Flagger
In this section, we discuss the stability of MMD-Flagger in detecting hallucinations, focusing on the impact of randomness and sample size. The detection process involves stochastic sampling, where a random selection of the next token is made based on the probability distribution generated by the softmax function, as detailed in Section 2.1. Our empirical findings suggest that MMD-Flagger is effective in identifying hallucinations with sequences comprising just $25$ tokens.
@@ Appendix F Detection Stability of MMD-Hallucination-Flagger
The MMD distance $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ plays a significant role in shaping the U-curve of $\mathcal{T}_{\widehat{MMD}}$. Therefore, we analyze the distribution of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ across 294 examples.
Our assumption is that if $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ is close to the converged value, the distance value does not change significantly, regardless of the random seeds used or the sample size. As a result, the distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ are expected to exhibit similar shapes.
Figure 7 illustrates the distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ at the first and third repetitions. With a sample size of 10, the two distributions appear quite different; however, at a sample size of 100, the distribution shapes seem to converge.
@@
Figure 7: Comparisons of distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ over $294$ examples. The left column shows the distributions at first repetition, and the right column is for the one of the third repetition.
Appendix G MMD trajectory smoothing
Table 13: Comparing the hallucination detection of the original MMD trajectory and smoothed trajectories. The score formats are “recall (precision)”.
Vector representation Kernel Smoothing Window size LFAN-HALL
avg dot No smoothing 0.820 (0.125)
@@ Appendix G MMD trajectory smoothing
This section introduces the hallucination detection performance after smoothing the MMD trajectory. Figure 8 plots an example of the smoothing, where the blue solid line is the original MMD trajectory, and purple and green dashed lines are smoothed lines. After the smoothing, a temperature $\tau$ at the argmin value change; the red point on the original MMD trajectory, while “x” marks on the smoothed lines. Even applying the smoothing, we do not observe substantial differences, as shown in Table 13.
The MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ can exhibit irregular shapes, as illustrated in Figure 8 and smoothed trajectories by the average window. To address these irregularities, a smoothing method can be applied to create a more uniform trajectory. By utilizing smoothing techniques, the MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ becomes easier to manage, which may enhance the ability of MMD-Flagger to detect hallucinations more effectively.
Table 13 compares the hallucination detection performance of the original MMD trajectory and smoothed trajectories. The length scale computation at $25$-th percentile; same as Section 3.2.
Figure 8: An example of the MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ and smoothed trajectories with the average window by $2$ and 33 sizes. The blue solid line is the original MMD trajectory. The purple dashed line is the smoothed line with the window size of 22, and the green dashed line is the smoothed line with the window size of $3$. The red circle is the minimum value of the original MMD trajectory, and black “x” marks are the minimum values on the smoothed trajectory.
Appendix H Qualitative results
The detection examples are all with the MMD-Flagger with the vector representation of avg and the kernel setting of gaussian using the $25$-th percentile value for computing the Gaussian kernel length scale, reported in Table 2. We note that the vertical axes in the following Figures are $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H_{\rm sto}^{\tau})$, which represents that we use the beam search for the decoding method at the next token prediction.
Correct hallucination detections (True Positive). Figures 9, 10, 11 illustrate a true positive detection, where the MMD-Flagger successfully identified a hallucinated translation ($y_{\text{hyp}}$ and $y_{\text{sto}}^{\tau=0.1}$) that significantly deviates from the source and reference texts.
$y_{\text{hyp}}$ The room was very small and the bathroom was very small.
$y_{\text{sto}}^{\tau=0.1}$ The hotel is located in a quiet area of the city.
Figure 9: True positive Example. The hypothesis output $y_{\rm hyp}$ is hallucinated. Descriptions are found in Appendix H.
$y_{\text{hyp}}$ Old and young people will find themselves together in the village’s respective squares for fun and fun.
Figure 10: True positive Example. The hypothesis output $y_{\rm hyp}$ is hallucinated. Descriptions are found in Appendix H.
$y_{\text{hyp}}$ The hotel is located in the heart of the city, just a few steps away from the train station.
Figure 11: True positive Example. The hypothesis output $y_{\rm hyp}$ is hallucinated. Descriptions are found in Appendix H.
Correct translations (True Negative). Figures 12, 13 show true negative cases. As these generated outputs $y_{\rm hyp}$ are correct, the MMD trajectory is the shape of monotonously increasing, and MMD-Flagger does not flag to these hypothesis outputs.
Reference The president concluded the debate as follows:
$y_{\text{hyp}}$ At the end of the debate, the President said:
Figure 12: True negative example. The hypothesis output $y_{\rm hyp}$ is correct output. Descriptions are found in Appendix H.
Source Politiken, auf die besonderes Augenmerk zu richten ist, und
Reference policies which should be given special attention, and
$y_{\text{hyp}}$ policies to which particular attention needs to be paid; and
Figure 13: True negative example. The hypothesis output $y_{\rm hyp}$ is correct output. Descriptions are found in Appendix H.
Detection by error (False Positive). At Figures 14, 15, the false positive cases are often observed when $y_{\text{hyp}}$ and $y_{\text{sto}}^{\tau=0.1}$ are semantically similar; however, dissimilarity due to different expressions or words. These false-positive errors would be due to the vector representations avg of the word embedding vectors. Due to the aggregation, the Gaussian kernel failed to compute the similarity correctly, and thus, $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ becomes a high value.
Source Ich komme zum Schluß noch darauf zu sprechen.
Reference I shall come back to this point again at the end.
$y_{\text{hyp}}$ Finally, I would like to mention this.
$y_{\text{sto}}^{\tau=0.1}$ I will end by mentioning this.
Figure 14: False positive example. The hypothesis output $y_{\rm hyp}$ is correct output, however, MMD-Flagger flags it as the hallucination. Descriptions are found in Appendix H.
Source Wollen sie eine Firma gründen oder umstrukturieren?
Reference Do you intend to market new product or service?
$y_{\text{hyp}}$ Do they want to set up or restructure a company?
$y_{\text{sto}}^{\tau=0.1}$ Do you want to start a company or restructure it?
Figure 15: False positive example. The hypothesis output $y_{\rm hyp}$ is correct output, however, MMD-Flagger flags it as the hallucination. Descriptions are found in Appendix H.
Fail to detect (False Negative). At Figures 16, 17, 18, the false negative errors are often observed when the stochastic sampled translations $y_{\text{sto}}^{\tau=0.1}$ are really similar to the beam search translation $y_{\text{hyp}}$. In this case, $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ becomes a small value and is hard to form a U-shape; therefore, MMD-Flagger failed to detect the hallucination.
Source Februar 1997 angenommenen Entschließung ABl.
Reference 61. adopted on 19 February 1997;
$y_{\text{hyp}}$ Resolution OJ L 208, p.
$y_{\text{sto}}^{\tau=0.1}$ Resolution OJ L 208, p.
Figure 16: False negative example. The hypothesis output $y_{\rm hyp}$ is hallucinated, however, MMD-Flagger fails to flag it as the hallucination. Descriptions are found in Appendix H.
Source Oktober zu verlängern.
Reference police and penitentiary.
$y_{\text{hyp}}$ The Commission has decided to extend the programme.
Figure 17: False negative example The hypothesis output $y_{\rm hyp}$ is hallucinated, however, MMD-Flagger fails to flag it as the hallucination. Descriptions are found in Appendix H.
$y_{\text{hyp}}$ Mr Martin BARTENSTEINBundesminister für Wirtschaft und Arbeit.
Figure 18: False negative example. The hypothesis output $y_{\rm hyp}$ is hallucinated, however, MMD-Flagger fails to flag it as the hallucination. Descriptions are found in Appendix H.